OpenStudy (anonymous):

A photon with a wavelength (λ) of 3.091 x 10–7 m strikes an atom of hydrogen. Determine the velocity (in meters/second) of an electron ejected from the excited state, n = 3.

OpenStudy (anonymous):

Using this equation: \[E_{incident photon}=E_{binding}+E_{scattered electron}\] And knowing: \[E_{incident}=-h*c/\lambda\] \[E_{scattered}=1/2*m*v^2\] \[E_{binding}=-K*(1/n^2)\] Being K=R*c*h (Rydberg constant, speed of light and Planck’s constant respectivily) Then you can get the velocity substituting.

OpenStudy (anonymous):

963943.3662